if a+b = 35
and 2a + 4b = 94
how do you find a and b?

Answer by boilpoil(127)

a+b=35

2a+4b=94 ---You need to make b the subject to insert the equation into 'b'
4b=94-2a
b=(94-2a)/4

a+(94-2a)/4=35
4a+94-2a=140
2a=46
a=23

b=(94-2(23))/4 ---Substitute a=23 into 'a'
b=(94-46)/4
b=48/4
b=12

And 23+12=35